Three +0.12C charges from an equilateral triangle 1.7m on a side. Using energy supplied


At the rate of 0.83KW, how many days would be required to move one of the charges to


the midpoint of the line joining the other two charges?

1
Expert's answer
2021-11-15T10:33:32-0500

The potential at the apex due to the other two charges is"k\times \frac{q1}{r} + k\times \frac{q2}{r}


= 9.0\times 10^9 \times(\frac{0.12}{1.70} + \frac{0.12}{1.70}) = 1.27\times10^9 V"


The potential between them is "9.0\times10^9 \times(\frac{0.12}{0.85} + \frac{0.12}{0.85}) = 2.54\times10^9 V"


so the potential difference is "2.54\times10^9 - 1.27\times10 9 = 1.27\times10^9 V"


so the energy "= V\times q = 1.27\times10^9 \times0.12 = 1.52\times10^8 J"

Since P ="\frac{Energy}{t}" then t "= \frac{E}{P} = \frac{1.52\times10^8}{0.83\times10^3} = 1.831\times10^5 s"


converting to days (1 day = 86400s) we get t ="\frac{ 1.831\times10^5}{86400} = 2.11" days